OpenStudy (love_to_love_you):

A jet ski company charges a flat fee of $26.00 plus $3.25 per hour to rent a jet ski. Another company charges a fee of $24.00 plus $3.50 per hour to rent the same jet ski. Find the number of hours for which the costs are the same. Round your answer to the nearest whole hour. a 2 b 8 c 10 d 12

OpenStudy (love_to_love_you):

@iGreen

OpenStudy (love_to_love_you):

@TuringTest

OpenStudy (love_to_love_you):

@amistre64

OpenStudy (amistre64):

so each company charges a flat fee and a variable fee. you will want to develop some line equations to play with

OpenStudy (love_to_love_you):

how?

OpenStudy (love_to_love_you):

@satellite73

OpenStudy (amistre64):

well, if we have to pay a flat fee, plus some fee based on an hourly rate .... how does this relate to a line equation?

OpenStudy (love_to_love_you):

26x+3.25y?

OpenStudy (amistre64):

i was thinking more of some y = mx + c variant :)

OpenStudy (amistre64):

costA = $26.00 + $3.25 per hour

OpenStudy (love_to_love_you):

y=26+3.25x?

OpenStudy (amistre64):

costB = $24.00 plus $3.50 per hour

OpenStudy (amistre64):

yes

OpenStudy (love_to_love_you):

ok...

OpenStudy (amistre64):

and the solution is where the lines meet

OpenStudy (love_to_love_you):

how do I find when the costs are the same?

OpenStudy (amistre64):

well, 2 lines cross where they have a pointin common

OpenStudy (amistre64):

costA = $26.00 + $3.25 per hour costB = $24.00 + $3.50 per hour when does A = B ? $26.00 + $3.25 per hour = $24.00 + $3.50 per hour solve for 'x'

OpenStudy (love_to_love_you):

where is x?

OpenStudy (amistre64):

x is a variable ... hour is variable

OpenStudy (amistre64):

..read between the lines ... between ,,, the lions. ah weema wah ah wee mah wah

OpenStudy (love_to_love_you):

what in the world haha

OpenStudy (amistre64):

got side tracked lol

OpenStudy (amistre64):

26.00 + 3.25x = 24 + $3.50x

OpenStudy (love_to_love_you):

50=6.75x 7.4=x

OpenStudy (love_to_love_you):

8? B?